OpenStudy (anonymous):

(X-27)/[(X^1/3)-3] find the limit as X approaches 27

OpenStudy (zarkon):

use\[x^3-a^3=(x-a)(x^2+xa+a^2)\]

OpenStudy (zarkon):

replace \[x\] with \[x^{1/3}\]

OpenStudy (anonymous):

and what about a?

OpenStudy (zarkon):

a=3

OpenStudy (zarkon):

\[x-27=(x^{1/3}-3)(x^{2/3}+x^{1/3}3+9)\]

OpenStudy (anonymous):

the first bracket simplifies to zero no?

OpenStudy (zarkon):

\[\frac{x-27}{x^{1/3}-3}=\frac{(x^{1/3}-3)(x^{2/3}+x^{1/3}3+9)}{x^{1/3}-3}\] \[=\frac{\cancel{(x^{1/3}-3)}(x^{2/3}+x^{1/3}3+9)}{\cancel{x^{1/3}-3}}=x^{2/3}+x^{1/3}3+9\] Now take limit...

OpenStudy (anonymous):